# Achieving Unbiased Multi-Instance Learning via Balanced Fine-Grained Positive-Unlabeled Learning

### v2

@@ C.1. Experiments with Low Imbalance Ratio
BFGPU $88.12\pm 1.24$ $88.36\pm 0.58$ $86.57\pm 0.69$ $84.61\pm 0.43$ $86.92$
Table 5. The table presents the F1 scores of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC with respect to changes in F1 on the SST-2 Dataset.
Methods F1 Score $AUC_{F1}$
2 4 6 8 10
Macro PN $83.63\pm 0.21$ $79.35\pm 1.04$ $78.15\pm 2.81$ $75.77\pm 5.79$ $75.50\pm 3.23$ $78.48$
MIL ATT $80.54\pm 0.51$ $62.39\pm 7.22$ $45.00\pm 36.50$ $00.00\pm 0$ 0.00 $48.79\pm 30.96$ $47.34$
FGSA $49.38\pm 43.89$ $44.44\pm 38.49$ $66.67\pm 0.00$ $44.44\pm 38.49$ $44.44\pm 38.49$ $49.87$
Micro DeepSAD $51.95\pm 12.30$ $14.95\pm 2.18$ $9.01\pm 7.18$ $9.92\pm 2.40$ $14.18\pm 4.29$ $20.00$
uPU $76.82\pm 3.40$ $64.43\pm 6.62$ $62.13\pm 6.88$ $60.79\pm 10.15$ $67.56\pm 1.27$ $66.35$
nnPU $84.00\pm 1.25$ $66.74\pm 0.10$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $70.15$
Table 6. The table presents AvgAcc and F1 scores of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$, along with the estimated AUC with respect to changes in AvgAcc and F1 on the Amazon Dataset.
@@
BFGPU $90.03\pm 0.08$ $88.51\pm 1.05$ $87.44\pm 0.17$ $86.50\pm 0.40$ $88.12$
Table 7. The table presents AvgAcc and F1 scores of different algorithms under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC with respect to changes in AvgAcc and F1 on the Sentiment140 Dataset.
@@
Macro PN $74.37\pm 0.40$ $69.80\pm 0.37$ $66.81\pm 0.48$ $65.09\pm 0.19$ $62.20\pm 0.48$ $67.65$
MIL ATT $71.27\pm 0.36$ $64.97\pm 3.24$ $52.88\pm 4.98$ $50.06\pm 0.10$ $50.16\pm 0.28$ $57.87$
FGSA $64.17\pm 3.13$ $53.46\pm 4.03$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $53.53$
PReNET $70.41\pm 3.74$ $57.11\pm 7.11$ $55.56\pm 4.83$ $54.85\pm 2.93$ $50.00\pm 0.00$ $57.59$
Micro DeepSAD $51.07\pm 2.14$ $51.80\pm 2.05$ $65.23\pm 2.04$ $63.99\pm 8.28$ $58.81\pm 8.64$ $58.18$
nnPU $69.66\pm 1.01$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $53.93$
robustPU $74.65\pm 0.70$ $70.13\pm 1.63$ $66.94\pm 0.40$ $59.14\pm 8.58$ $62.88\pm 0.27$ $66.74$
BFGPU $75.86\pm 1.01$ $70.68\pm 0.97$ $67.95\pm 1.53$ $66.34\pm 1.25$ $64.45\pm 0.35$ $69.06$
@@
Macro PN $73.99\pm 0.72$ $68.75\pm 0.44$ $64.94\pm 1.66$ $63.86\pm 0.52$ $59.50\pm 1.48$ $66.21$
MIL ATT $67.64\pm 0.52$ $56.56\pm 5.58$ $37.40\pm 34.07$ $22.85\pm 37.96$ $23.34\pm 37.56$ $41.56$
FGSA $52.19\pm 8.97$ $16.28\pm 17.59$ $44.44\pm 38.49$ $44.44\pm 38.49$ $44.44\pm 38.49$ $40.36$
Micro DeepSAD $54.65\pm 1.80$ $55.89\pm 1.76$ $67.50\pm 1.80$ $66.55\pm 7.42$ $61.86\pm 7.61$ $61.29$
uPU $60.88\pm 2.42$ $57.32\pm 10.29$ $48.65\pm 13.84$ $60.66\pm 4.54$ $66.66\pm 0.21$ $58.83$
nnPU $75.50\pm 0.56$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $68.44$
BFGPU $77.21\pm 0.73$ $73.27\pm 0.95$ $69.80\pm 3.18$ $67.21\pm 2.94$ $66.40\pm 0.68$ $\bm{70.78}$
Table 8. Experimental Results on the SST-2 Dataset with Imbalance at Both Macro and Micro Levels.
@@
Macro PN $72.65\pm 2.15$ $70.46\pm 4.82$ $69.32\pm 1.90$ $63.52\pm 6.64$ $55.42\pm 3.59$ $66.27$
DeepSAD $51.93\pm 1.63$ $54.79\pm 2.63$ $52.17\pm 1.77$ $45.91\pm 2.71$ $49.22\pm 1.98$ $50.80$
MIL ATT $77.90\pm 2.91$ $71.29\pm 14.46$ $51.21\pm 1.51$ $56.29\pm 7.21$ $58.14\pm 14.10$ $62.96$
FGSA $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00$
PReNET $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00$
Micro DeepSAD $50.37\pm 2.14$ $53.13\pm 2.07$ $49.52\pm 2.80$ $45.60\pm 3.80$ $50.39\pm 3.95$ $49.80$
DeepSVDD $51.38\pm 1.99$ $54.29\pm 2.23$ $50.00\pm 6.32$ $48.74\pm 3.93$ $53.10\pm 6.40$ $51.50$
uPU $71.45\pm 15.17$ $77.89\pm 1.23$ $57.25\pm 7.49$ $50.00\pm 0.00$ $50.00\pm 0.00$ $61.32$
nnPU $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00$
balancedPU $54.24\pm 7.34$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.00\pm 0.00$ $50.85$
robustPU $75.60\pm 5.76$ $76.57\pm 2.73$ $58.21\pm 14.22$ $50.00\pm 0.00$ $50.00\pm 0.00$ $62.08$
BFGPU $80.48\pm 3.23$ $78.22\pm 2.48$ $73.19\pm 4.41$ $63.84\pm 11.18$ $67.83\pm 0.46$ $72.71$
@@
Macro PN $78.16\pm 1.37$ $76.95\pm 2.89$ $75.92\pm 1.08$ $72.56\pm 3.22$ $68.00\pm 1.67$ $74.32$
DeepSAD $56.88\pm 3.37$ $59.33\pm 5.46$ $44.46\pm 6.05$ $36.40\pm 14.12$ $23.49\pm 14.07$ $44.11$
MIL ATT $81.46\pm 1.93$ $77.75\pm 8.28$ $67.21\pm 0.68$ $69.70\pm 3.55$ $70.73\pm 7.03$ $73.37$
FGSA $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67$
PReNET $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67$
Micro DeepSAD $47.48\pm 3.96$ $51.85\pm 1.29$ $44.26\pm 3.04$ $38.50\pm 2.66$ $47.01\pm 5.18$ $45.82$
DeepSVDD $48.99\pm 4.43$ $51.81\pm 2.14$ $44.85\pm 6.23$ $48.33\pm 11.22$ $50.36\pm 4.26$ $48.87$
uPU $77.09\pm 7.37$ $78.81\pm 0.83$ $68.25\pm 2.24$ $66.67\pm 0.00$ $66.67\pm 0.00$ $71.50$
nnPU $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67\pm 0.00$ $66.67$
balancedPU $13.85\pm 23.99$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $2.77$
robustPU $80.07\pm 3.93$ $80.74\pm 1.57$ $71.03\pm 7.56$ $66.67\pm 0.00$ $66.67\pm 0.00$ $73.04$
BFGPU $83.48\pm 2.32$ $81.59\pm 1.29$ $78.76\pm 2.76$ $64.92\pm 18.60$ $\bm{74.24\pm 3.95}$ $76.60$
Table 9. The table presents AvgAcc and F1 results of the ablation study under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$ on the IMDB dataset.
@@
$\checkmark$ $\checkmark$ $\checkmark$ $88.12\pm 1.24$ $88.36\pm 0.58$ $86.57\pm 0.69$ $84.61\pm 0.43$
Table 10. The table presents AvgAcc and F1 results of the ablation study under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$ on the SST-2 dataset.
@@
$2$ $4$ $6$ $8$ $10$
$\times$ $\checkmark$ $\checkmark$ $85.27\pm 1.50$ $83.33\pm 0.84$ $79.95\pm 2.46$ $77.36\pm 2.67$ $79.07\pm 3.80$
@@
Table 11. This table presents the AvgAcc of the sensitive analysis of $\lambda_{bfgpu}$ on the SST-2 dataset.
@@
Table 12. This table presents the AvgAcc of the sensitive analysis of $\lambda_{pse}$ on the SST-2 dataset.
@@
C.2. Experiments with High Imbalance Ratio
Due to space limitations in the main text, the additional experiments on the Sentiment140 datasets are provided in table 7 in this appendix.
C.3. Experiments with Imbalance at Both Macro and Micro Levels
On the SST-2 dataset, we set $\sigma_{macro}$ to 5 and explored how algorithm performance varied with $\sigma_{micro}$ set to [2, 4, 6, 8, 10].
Due to space limitations in the main text, the experimental statistical data on the SST-2 dataset are provided in table 8 in this appendix.
C.4. Ablation Study
To explore the necessity of each component of BFGPU, we conducted ablation experiments, comparing the algorithm’s performance without the new loss function, without pseudo-labeling, and without the threshold adjustment technique. We performed these experiments on the IMDB and SST-2 datasets, setting $\sigma_{micro}$ to [2, 4, 6, 8, 10]. Due to space limitations in the main text, the additional experiments on the SST-2 dataset are provided in table 10 in this appendix.
C.5. Sensitive Analysis
We have supplemented the sensitivity analysis of hyperparameters $\lambda_{bfgpu}$ and $\lambda_{pse}$. We completed experiments on the IMDB dataset, setting $\sigma_{micro}=5$. We set the variation range of the two parameters to [1, 2, 3, 4, 5], and kept one constant while varying the other. Due to space limitations in the main text, the sensitive analysis of $\lambda_{pse}$ on the SST-2 dataset are provided in table 12 in this appendix.

### v3

@@ D.1 Experiments with Synthetic Datasets
BFGPU $8.12\pm 1.24 88$$8.36\pm 0.58$ $86.57\pm 0.69$ $84.61\pm 0.43$ $86.92$
Table 8: The table presents AvgAcc and F1 scores of various algorithms under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$, along with the estimated AUC with respect to changes in AvgAcc and F1 on the Amazon Dataset.
@@
BFGPU $90.03\pm 0.08$ $88.51\pm 1.05$ $87.44\pm 0.17$ $6.50\pm 0.40 88$.1288.12
D.2 Experiments with Imbalance at Both Macro and Micro Levels
On the SST-2 dataset, we set $\sigma_{macro}$ to 5 and 10 and explored how algorithm performance varied with $\sigma_{micro}$ set to [2, 4, 6, 8, 10].
Due to space limitations in the main text, the experimental statistical data on the SST-2 dataset are provided in tables˜9 and 10 in this appendix.
Table 9: Experimental Results with Imbalance at Both Macro and Micro Levels when $\sigma_{macro=5}$.
@@ D.2 Experiments with Imbalance at Both Macro and Micro Levels
Macro Supervised $72.65$ $70.46$ $69.32$ $63.52$ $55.42$ $66.27$
DeepSAD $51.93$ $54.79$ $52.17$ $45.91$ $49.22$ $50.80$
UnderSampling $75.41$ $73.27$ $62.32$ $55.81$ $54.64$ $64.29$
OverSampling $78.18$ $73.76$ $69.57$ $60.47$ $59.67$ $68.33$
MIL MaxPooling $50.00$ $50.00$ $50.00$ $50.00$ $57.87$ $51.57$
TopkPooling - $74.64$ $69.31$ $58.49$ $50.00$ $63.11$
Attention $77.90$ $71.29$ $51.21$ $56.29$ $58.14$ $62.96$
FGSA $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
PReNET $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
Micro DeepSAD $50.37$ $53.13$ $49.52$ $45.60$ $50.39$ $49.80$
DeepSVDD $51.38$ $54.29$ $50.00$ $48.74$ $53.10$ $51.50$
uPU $71.45$ $77.89$ $57.25$ $50.00$ $50.00$ $61.32$
nnPU $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
balancedPU $54.24$ $50.00$ $50.00$ $50.00$ $50.00$ $50.85$
robustPU $75.60$ $76.57$ $58.21$ $50.00$ $50.00$ $62.08$
BFGPU $80.48$ $78.22$ $73.19$ $63.84$ $67.83$ $72.71$
@@
Macro Supervised $78.16$ $76.95$ $75.92$ $72.56$ $68.00$ $74.32$
DeepSAD $56.88$ $59.33$ $44.46$ $36.40$ $23.49$ $44.11$
UnderSampling $69.83$ $75.00$ $71.28$ $69.84$ $56.86$ $68.56$
OverSampling $70.56$ $65.79$ $62.37$ $60.47$ $62.90$ $64.42$
MIL MaxPooling $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
TopkPooling - $75.00$ $69.38$ $68.00$ $66.67$ $69.76$
Attention $81.46$ $77.75$ $67.21$ $69.70$ $70.73$ $73.37$
FGSA $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
PReNET $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
Micro DeepSAD $47.48$ $51.85$ $44.26$ $38.50$ $47.01$ $45.82$
DeepSVDD $48.99$ $51.81$ $44.85$ $48.33$ $50.36$ $48.87$
uPU $77.09$ $78.81$ $68.25$ $66.67$ $66.67$ $71.50$
nnPU $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
balancedPU $13.85$ $0.00$ $0.00$ $0.00$ $0.00$ $2.77$
robustPU $80.07$ $80.74$ $71.03$ $66.67$ $66.67$ $73.04$
BFGPU $83.48$ $81.59$ $78.76$ $64.92$ $\bm{74.24}$ $76.60$
Table 10: Experimental Results with Imbalance at Both Macro and Micro Levels when $\sigma_{macro=10}$.
@@
Macro Supervised 68.78 54.95 50.00 50.00 50.00 54.75
DeepSAD 66.02 56.93 53.62 39.06 55.81 54.29
UnderSampling 77.62 70.79 60.14 50.00 53.48 62.41
OverSampling 69.89 66.83 61.59 62.26 54.65 63.04
MIL MaxPooling $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
TopkPooling - 72.77 50.72 50.00 50.00 55.87
Attention 75.41 50.99 71.01 50.00 50.00 59.48
FGSA $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
PReNET 77.62 72.77 65.94 74.52 50.00 68.17
Micro DeepSAD 47.51 51.49 43.48 49.06 47.67 47.84
DeepSVDD 56.60 58.49 51.16 58.14 45.35 53.95
uPU 74.03 66.34 50.00 50.00 50.00 58.07
nnPU $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
balancedPU $50.00$ $50.00$ $50.00$ $50.00$ $50.00$ $50.00$
robustPU 74.86 63.37 50.00 50.00 50.00 57.65
BFGPU $79.01$ $78.22$ $78.99$ $76.42$ $68.60$ $76.25$
@@
Macro Supervised 75.80 68.51 66.67 66.67 66.67 68.86
DeepSAD 74.43 65.32 53.62 56.88 32.14 56.48
UnderSampling 77.56 63.80 49.54 66.67 23.08 56.13
OverSampling 78.87 70.73 74.36 76.19 66.67 73.36
MIL MaxPooling 0.00 0.00 0.00 0.00 0.00 0.00
TopkPooling - 78.43 66.99 66.67 66.67 69.69
Attention 79.91 67.11 77.27 66.67 66.67 71.53
FGSA $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
PReNET 81.29 79.39 78.26 74.59 66.67 76.04
Micro DeepSAD 53.66 59.17 50.63 51.79 60.18 55.09
DeepSVDD 64.26 44.75 59.63 63.83 67.19 59.93
uPU 83.56 71.67 $66.67$ $66.67$ $66.67$ 71.05
nnPU $66.67$ $66.67$ $66.67$ $66.67$ $66.67$ $66.67$
balancedPU 0.00 0.00 0.00 0.00 0.00 0.00
robustPU 79.55 73.19 66.67 66.67 66.67 70.55
BFGPU $87.27$ $81.48$ $78.52$ $76.19$ $67.19$ $78.13$
D.3 Ablation Study
To explore the necessity of each component of BFGPU, we conducted ablation experiments, comparing the algorithm’s performance without the new loss function, without pseudo-labeling, and without the threshold adjustment technique. We performed these experiments on the IMDB and SST-2 datasets, setting $\sigma_{micro}$ to [2, 4, 6, 8, 10]. Due to space limitations in the main text, the additional experiments on the SST-2 dataset are provided in table˜12 in this appendix.
Table 11: The table presents AvgAcc and F1 results of the ablation study under varying values of $\sigma_{\text{micro}}\in\{2,3,4,5\}$ on the IMDB dataset.
@@ D.3 Ablation Study
$\checkmark$ $\checkmark$ $\checkmark$ $8.12\pm 1.24 88$$8.36\pm 0.58$ $86.57\pm 0.69$ $84.61\pm 0.43$
Table 12: The table presents AvgAcc and F1 results of the ablation study under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$ on the SST-2 dataset.
@@
22 44 66 88 $10$
$\times$ $\checkmark$ $\checkmark$ $85.27\pm 1.50$ $83.33\pm 0.84$ $9.95\pm 2.46 77$$7.36\pm 2.67$ $79.07\pm 3.80$
@@
D.4 Experiments with Different Base Model
To verify that our results are not limited by the choice of backbone model, we conducted further evaluations using DeBERTa. The results are shown in table˜13.
Table 13: The table presents AvgAcc and F1 scores of various algorithms with the Deberta model under varying values of $\sigma_{\text{micro}}\in\{2,4,6,8,10\}$, along with the estimated AUC with respect to changes in AvgAcc and F1 on the SST-2 Dataset.
Method AvgAcc $AUC_{AvgAcc}$
2 4 6 8 10
Macro Supervised 81.22 77.72 78.26 77.36 77.91 78.49
DeepSAD 49.17 59.41 42.75 50.94 47.67 49.99
MIL MaxPooling 75.97 70.30 76.81 65.09 50.00 67.63
TopkPooling - 74.26 77.54 78.30 73.26 75.84
Attention 72.65 65.35 52.90 50.83 50.94 58.53
FGSA 73.48 64.36 54.35 50.94 50.00 58.63
PReNET 71.82 61.39 65.22 55.66 61.63 63.14
Micro DeepSAD 48.34 49.51 51.45 51.89 52.33 50.70
DeepSVDD 48.07 40.10 52.90 51.89 51.16 48.82
uPU 72.65 57.43 72.46 59.43 67.44 65.88
nnPU 72.65 50.50 50.00 50.00 50.00 54.63
balancedPU 82.32 80.20 78.26 76.42 72.93 78.03
robustPU 83.98 80.20 76.47 83.33 80.19 80.50
BFGPU $88.40$ $81.68$ $78.99$ $83.96$ $81.40$ $82.89$
Method F1 Score $AUC_{F1}$
2 4 6 8 10
Macro Supervised 80.79 77.61 79.73 76.92 79.12 78.83
DeepSAD 55.34 69.40 56.83 60.61 62.81 61.00
MIL MaxPooling 65.37 63.42 59.76 50.04 53.70 58.46
TopkPooling - 63.23 64.87 53.44 55.46 59.00
FGSA 65.71 55.56 22.22 3.70 0.00 29.44
Attention 63.47 47.76 35.29 23.08 10.71 36.06
PReNET 74.75 40.00 47.83 25.40 37.74 45.14
Micro DeepSAD 63.55 66.23 57.32 61.65 67.72 63.29
DeepSVDD 64.26 44.75 59.63 63.83 67.19 59.93
uPU 64.52 33.85 66.67 41.10 62.16 53.66
nnPU 78.15 66.89 66.67 66.67 66.67 69.01
balancedPU 84.08 82.91 81.71 79.67 81.19 81.91
robustPU 82.53 77.78 80.77 82.44 75.32 79.77
BFGPU $88.65$ $84.12$ $81.76$ $85.96$ $83.33$ $84.76$
D.5 Experiments with Different PU Loss with Pseudo Labeling and ADT
o further demonstrate the superiority of the BFGPU loss function, we compared its performance against other PU learning methods when integrated with our proposed pseudo-labeling and ADT strategies. The results are shown in table˜14.
Table 14: The table presents the performance of various PU learning methods when integrated with our proposed pseudo-labeling and ADT strategies.
Method AvgAcc $AUC_{AvgAcc}$
2 4 6 8 10
uPU $82.69\pm 2.02$ $76.07\pm 1.82$ $71.26\pm 3.36$ $64.47\pm 0.89$ $61.24\pm 4.78$ $71.15$
nnPU $87.02\pm 0.60$ $83.83\pm 1.42$ $79.23\pm 1.49$ $75.16\pm 2.48$ $74.03\pm 3.95$ $79.85$
balancedPU $9.83\pm 6.66 77$$7.23\pm 5.73$ $8.84\pm 11.79 66$$6.04\pm 5.82$ $72.09\pm 1.64$ $72.81$
robustPU $86.00\pm 1.84$ $74.09\pm 20.01$ $78.02\pm 2.33$ $72.96\pm 8.02$ $61.63\pm 14.52$ $74.54$
Method F1 Scroe $AUC_{F1}$
2 4 6 8 10
uPU $84.09\pm 1.08$ $77.99\pm 1.55$ $75.59\pm 2.66$ $69.91\pm 1.50$ $65.42\pm 9.19$ $74.60$
nnPU $87.75\pm 0.60$ $85.13\pm 0.84$ $81.56\pm 1.00$ $78.85\pm 2.79$ $78.39\pm 2.57$ $82.34$
balancedPU $80.94\pm 6.40$ $79.54\pm 4.81$ $57.02\pm 32.67$ $65.03\pm 13.16$ $76.11\pm 1.94$ $71.73$
robustPU $86.91\pm 1.48$ $59.86\pm 45.37$ $80.23\pm 2.51$ $74.05\pm 8.97$ $48.39\pm 42.41$ 69.89
D.6 Sensitive Analysis
We have supplemented the sensitivity analysis of hyperparameters $\lambda_{bfgpu}$ and $\lambda_{pse}$. We completed experiments on the IMDB dataset, setting $\sigma_{micro}=5$. We set the variation range of the two parameters to [1, 2, 3, 4, 5], and kept one constant while varying the other. Due to space limitations in the main text, the sensitive analysis of $\lambda_{pse}$ on the SST-2 dataset are provided in table˜16 in this appendix.
Table 15: This table presents the AvgAcc of the sensitive analysis of $\lambda_{bfgpu}$ on the SST-2 dataset.
@@ D.6 Sensitive Analysis
Table 16: This table presents the AvgAcc of the sensitive analysis of $\lambda_{pse}$ on the SST-2 dataset.